# Test-Time Scaling with World Models for Spatial Reasoning

### v1

@@ Abstract
Spatial reasoning in 3D space is central to human cognition and indispensable for embodied tasks such as navigation and manipulation. However, state-of-the-art vision–language models (VLMs) struggle frequently with tasks as simple as anticipating how a scene will look after an egocentric motion: they perceive 2D images but lack an internal model of 3D dynamics. We therefore propose MindJourney, a test-time scaling framework that grants a VLM with this missing capability by coupling it to a controllable world model based on video diffusion. The VLM iteratively sketches a concise camera trajectory, while the world model synthesizes the corresponding view at each step. The VLM then reasons over this multi-view evidence gathered during the interactive exploration. Without any fine-tuning, our MindJourney achieves over an average 8% performance boost on the representative spatial reasoning benchmark SAT, showing that pairing VLMs with world models for test-time scaling offers a simple, plug-and-play route to robust 3D reasoning. Meanwhile, our method also improves upon the test-time inference VLMs trained through reinforcement learning, which demonstrates the potential of our method that utilizes world models for test-time scaling.
@@ 1 Introduction
Recent advances in vision–language models (VLMs) have produced systems that are already comparable to human performance in visual recognition and comprehension, even solving simple spatial-reasoning problems (Chen et al., 2024a; Zhang et al., 2025b). However, these models still fall short of reasoning about the actual 3D world that lies behind a 2D image as humans do. Recent spatial reasoning benchmarks (Yang et al., 2024a; Ray et al., 2025; Zhang et al., 2025b, a) reveal that VLMs struggle on tasks that require imagining the consequences of egocentric movements. Given a question shown in the upper-left corner of Fig. 1, a state-of-the-art VLM easily fails on such a perspective-shift question, while given the same view, humans can solve it effortlessly by mentally simulating a short walk and visualizing the next view. This gap highlights a critical bottleneck: current VLMs do not yet treat an image as an interactive world, limiting their usefulness for embodied agents operating in 3D space.
@@
In a comprehensive evaluation on the SAT benchmark, SpatialNavigator achieves substantial gains on multiple spatial reasoning tasks. Our method improves top-1 accuracy performance across four very different VLM back-ends with two distinct world models by an average 8.1%, with the largest single gain reaching 15%. Our method also outperforms and can improve upon the o1-like test-time scaling VLMs, which demonstrates the compatibility of our method and showcases the potential of the idea of test-time scaling with world models.
@@ 3.2 World Model Formulation
Action-Driven Video Generation.
@@ 3.3 Spatial Beam Search for Action Space Exploration
[…]
@@
The global observation set for the current search step is the union of $\mathcal{O}(\tau)$ over all beam nodes. In practice, because the world model $\mathcal{W}$ renders an entire video clip in one pass, we roll out only the candidates in $\mathcal{C}(\tau)$ whose length is exactly $m+k$—that is, trajectories that extend the current node by the full $k$ actions. A single call to $\mathcal{W}$ with this batched set then produces the complete local observation set $\mathcal{O}(\tau)$ for the node. Fig. 3 illustrate this process with an example.
@@
$s_{\text{exp}}(\tau^{\prime})\quad\text{(``how useful is it to \emph{keep exploring} this trajectory?'')},$?”) ,
$s_{\text{help}}(\tau^{\prime})\quad\text{(``how useful is \emph{this view} foranswering now?'')}.$?”) .
@@ 3.4 Search World Model Details
Architecture. SWM is based on the CogvideoX-5b (Yang et al., 2024b), following CameraCtrl (He et al., 2024) and Memory-WM (Zhou et al., 2025b). Specifically, we represent the camera transform by Plücker embedding (Sitzmann et al., 2021) and directly concatenating them with input image channels for precise camera control. We extend the channel of the input layer in CogvideoX to accept additional conditions of camera embedding.
@@ 4.1 Experiment Settings
Vision-Language Models. We pair our pipeline with four representative VLMs: the closed-source GPT-4o and GPT-4.1 (strong general-purpose multimodal baselines), InternVL3-14B (one of the most capable open-source VLMs to date), and o1, a reinforcement-learning–fine-tuned model that performs test-time chain-of-thought scaling. This mix spans both proprietary and fully open ecosystems as well as models with and without explicit reasoning scaffolds. World Models. Our experiments use two distinct video world models. (i) Search World Model (SWM), the world model we trained, introduced in Sec. 3.4; and (ii) Stable-Virtual-Camera (SVC), a recently released, publicly available generator that produces geometrically stable novel views.
@@ 4.2 Experiments Results
The SAT-Real split comprises 150 real-image queries spanning indoor and outdoor scenes. Table 1 shows that augmenting each VLM with MindJourney results in uniform and sizeable gains. Average top-1 accuracy rises by 8.1 percentage points, and GPT-4.1 achieves the largest single boost at nearly 15%. Remarkably, GPT-4.1 with our method already surpasses vanilla o1; when o1 itself is paired with MindJourney it sets a new state of the art on SAT-Real. These findings confirm that world-model-driven test-time scaling complements RL-based scaling and our method generalizes to real, outdoor scenarios.
@@
Because the synthetic split contains 4 000 questions, we evaluate on a random 500-question subset to keep the o1 runs tractable. Results in Table 2 reproduce the pattern seen on real images: mean accuracy improves by 8.5% in average. Again, the performance of GPT-4.1 with our method surpasses vanilla o1, and o1 with our method improves its performance. Moreover, across all five SAT question types the highest score is always achieved by a MindJourney-augmented model, underscoring that the proposed framework consistently improves reasoning across all question types.
@@ 4.3 Ablation Study
In our ablation study, we ablate the hyperparameters of our search method. Specifically, for gpt-4o augmented with SWM, we ablate with search step $n\in\{1,2,3\}$ and VLM pruning threshold $\gamma\in\{5,7,9\}$. We experimented on these values on both SAT-Real and SAT-Synthesized. Accuracy is reported on both SAT-Real and the 500-question SAT-Synthesized subset, allowing us to evaluate how each setting affects performance in real-image and synthetic scenarios.
Search Depth. Panel (a) of Fig. 4 shows that accuracy on SAT-Real peaks at search steps 2 and then drops slightly at step 3. The decline stems from the limits of our world model, SWM: trained predominantly on indoor synthetic data, it struggles to faithfully simulate outdoor views once the imagined trajectory strays too far from the reference frame. In contrast, on SAT-Synthesized—whose scenes are closer to the SWM’s training distribution—performance continues to rise through step 3, confirming that deeper exploration remains beneficial when the world model can faithfully render long roll-outs. More visual results can be found in the Appendix.
Pruning Threshold. Panels (a) and (b) also highlight the crucial role of the VLM score threshold. A lenient threshold admits many low-quality views into the evidence buffer, diluting the signal and degrading final accuracy; an overly strict threshold, on the other hand, prematurely discards potentially informative trajectories.
@@
Figure 4: Inference Steps vs. Accuracy. Accuracy on SAT-Real and SAT-Synthesized with different VLM Thresholds.
@@ A.1.1 Search Configurations
We use the same search configuration for every experiment: a search depth of $n=3$ steps, a beam size of 2, exploration and helpfulness thresholds $\gamma_{\text{exp}}=8$ and $\gamma_{\text{help}}=8$, and a maximum trajectory length of 8 starting from the given reference image. During each expansion, we allow up to $k=3$ consecutive repetitions per primitive action; each forward step moves the agent $0.25$m and each rotation step turns it by $9^{\circ}$. We prepared more visual examples about the Trajectory Expansion process under our experiment settings at Fig. 5 and Fig. 6
@@ A.1.2 Computational Resources
All inference experiments were run on high-performance NVIDIA GPUs: when using Search World Model as the world model, we employed A40 GPUs with 40GB of VRAM; when using Stable-Virtual-Camera as the world model, we ran on H100 GPUs with 80GB of VRAM; and for all experiments combining the InternVL3-14B VLM with the Search World Model, we also used H100 GPUs to accommodate the larger memory footprint of the vision–language model.
@@ A.2.1 Dataset
The training set for our Search World Model (SWM) comprises three components: HM3D, DL3DV-10K, and RealEstate10K (Szot et al., 2022; Ling et al., 2024; Zhou et al., 2018).
HM3D. We generate 50K simulated navigation clips using Habitat on HM3D scenes. For each episode, we sample a random start and goal position and follow the shortest path for up to 500 steps. To improve robustness to camera tilt, we uniformly draw an initial pitch from $[-30^{\circ},20^{\circ}]$, and in 10% of episodes we hold the agent at a fixed location and vary only its pitch by rotating up and down.
@@
DL3DV-10K. Similarly, we incorporate 10K real outdoor videos from DL3DV-10K to capture natural scenes and broaden our model’s generalization to exterior settings.
@@ A.2.2 Implementation Details
Model Architecture. We adopt CogVideoX-5B (Yang et al., 2024b) [^2] https://github.com/THUDM/CogVideo; Apache-2.0 as our backbone. To incorporate camera embeddings, we expand the input channels of its PatchEmbed layer—duplicating the pretrained convolutional weights into the appropriate submatrix and zero-initializing all new parameters. Furthermore, we introduce LoRA adapters within each self-attention block, configured with rank $r=128$ and scaling factor $\alpha=128$, enabling parameter-efficient fine-tuning without disturbing the original transformer weights.
Action Representation. We represent each relative camera transform using the Plücker embedding (Sitzmann et al., 2021). Given intrinsic matrix $\mathbf{K}$ and extrinsic matrix $\mathbf{E}=[\mathbf{R}\;|\;\mathbf{t}]$, we build a 6-channel “image”
$P\in\mathbb{R}^{h\times w\times 6},$
where at pixel $(u,v)$ we store the pair
$\bigl{(}o_{(u,v)}\times d_{(u,v)},\,d_{(u,v)}\bigr{)}.$
Here
$o_{(u,v)}=\mathbf{t},\quad d_{(u,v)}=\mathbf{R}\,\mathbf{K}^{-1}\begin{pmatrix}u\\ v\\ 1\end{pmatrix}.$
This encoding lets the video model reason directly over 3D ray origins and directions for precise spatial control.
Decoupling Pitch. Our primitive actions $\{\mathtt{move\;forward},\;\mathtt{turn\;right},\;\mathtt{turn\;left}\}$ always preserve the camera’s pitch angle $\theta$. To enforce this, we factor each extrinsic transform $\mathbf{E}\in\mathrm{SE}(3)$ into a pure pitch rotation and a remaining horizontal motion:
$\mathbf{E}\;=\;R_{\mathrm{pitch}}(\theta)\;T_{\mathrm{horiz}},$
where $R_{\mathrm{pitch}}(\theta)$ is a rotation about the camera’s local $x$–axis:
$R_{\mathrm{pitch}}(\theta)=\begin{pmatrix}1&0&0&0\\ 0&\cos\theta&-\sin\theta&0\\ 0&\sin\theta&\cos\theta&0\\ 0&0&0&1\end{pmatrix},$
and the horizontal motion component is
$T_{\mathrm{horiz}}=R_{\mathrm{pitch}}(-\theta)\,\mathbf{E}.$
We then apply the Plücker embedding only to $T_{\mathrm{horiz}}$, producing a tensor in $\mathbb{R}^{h\times w\times 6}$. Finally, we broadcast the scalar pitch $\theta$ into an $(h\times w\times 1)$ map and concatenate it along the channel dimension to obtain the full $\mathbb{R}^{h\times w\times 7}$ action representation.
@@ A.2.3 Training Details
We subsample training clips with a frame-skip stride uniformly drawn between 1 and 3 to expose the model to varied camera motions. Optimization is performed with Adam and a linear warmup schedule to a peak learning rate of $41e-$ 4, using bfloat16 precision for efficiency and clipping gradients to a maximum norm of 1.0 for stability. All video-diffusion models are trained on eight NVIDIA H100 GPUs over approximately three days. At training time we adopt the v-prediction objective (Salimans and Ho, 2022) and, during inference, employ the DDIM sampler (Rombach et al., 2022) with 50 sampling steps—producing nine-frame videos in roughly nine seconds.
@@ B.1.1 Case 1: Inaccurate Forward Movement
In group (a), the imagined trajectories systematically under- or over-estimate forward translations: the actual step lengths no longer match the intended distances, and the displacement between successive frames varies unpredictably. As a result, the agent repeatedly overshoots its targets and exhibits jittery, erratic motion within the simulated environment.

### v2

@@ Abstract
Spatial reasoning in 3D space is central to human cognition and indispensable for embodied tasks such as navigation and manipulation. However, state-of-the-art vision–language models (VLMs) struggle frequently with tasks as simple as anticipating how a scene will look after an egocentric motion: they perceive 2D images but lack an internal model of 3D dynamics. We therefore propose MindJourney, a test-time scaling framework that grants a VLM with this missing capability by coupling it to a controllable world model based on video diffusion. The VLM iteratively sketches a concise camera trajectory, while the world model synthesizes the corresponding view at each step. The VLM then reasons over this multi-view evidence gathered during the interactive exploration. Without any fine-tuning, our MindJourney achieves an average 7.7% performance boost on the representative spatial reasoning benchmark SAT, showing that pairing VLMs with world models for test-time scaling offers a simple, plug-and-play route to robust 3D reasoning. Meanwhile, our method also improves upon the test-time inference VLMs trained through reinforcement learning, which demonstrates the potential of our method that utilizes world models for test-time scaling.
@@ 1 Introduction
Recent advances in vision–language models (VLMs) have produced systems that are already comparable to human performance in visual recognition and comprehension, even solving simple spatial-reasoning problems (Chen et al., 2024a; Zhang et al., 2025b). However, these models still fall short of reasoning about the actual 3D world that lies behind a 2D image as humans do. Recent spatial reasoning benchmarks (Yang et al., 2024; Ray et al., 2025; Zhang et al., 2025b, a) reveal that VLMs struggle on tasks that require imagining the consequences of egocentric movements. Given a question shown in the upper-left corner of Fig. 1, a state-of-the-art VLM easily fails on such a perspective-shift question, while given the same view, humans can solve it effortlessly by mentally simulating a short walk and visualizing the next view. This gap highlights a critical bottleneck: current VLMs do not yet treat an image as an interactive world, limiting their usefulness for embodied agents operating in 3D space.
@@
In a comprehensive evaluation on the SAT benchmark, SpatialNavigator achieves substantial gains on multiple spatial reasoning tasks. Our method improves top-1 accuracy performance across four very different VLM back-ends with two distinct world models by an average 7.7%, with the largest single gain over 10%. Our method also outperforms and can improve upon the o1-like test-time scaling VLMs, which demonstrates the compatibility of our method and showcases the potential of the idea of test-time scaling with world models.
@@ 3.2 World Model Formulation
Action‐Driven Video Generation.
@@ 3.3 Spatial Beam Search for Action Space Exploration
$\mathcal{C}(\tau)\;=\;\bigl\{\tau\!\oplus\!a^{\prime\,r}\;\big|\;a^{\prime}\in\mathcal{A},\;1\leq r\leq k\bigr\},\qquad|\mathcal{C}(\tau)|\leq 3k,$
@@
The global observation set for the current search step is the union of $\mathcal{O}(\tau)$ over all beam nodes. In practice, because the world model $\mathcal{W}$ renders an entire video clip in one pass, we roll out only the candidates in $\mathcal{C}(\tau)$ whose length is exactly m+km+k—that is, trajectories that extend the current node by the full $k$ actions. A single call to $\mathcal{W}$ with this batched set then produces the complete local observation set $\mathcal{O}(\tau)$ for the node. Fig. 3 illustrate this process with an example.
@@
[…]
$s_{\text{help}}(\tau^{\prime})\quad\text{(``how useful is \emph{this view} for answering now?'')}.$
@@ 3.4 Search World Model Details
Architecture. SWM is based on the Wan2.2-TI2V-5B (Wan et al., 2025), following ReCamMaster (Bai et al., 2025). Specifically, we represent the camera transform by camera extrinsic matrices and directly add the embeded camera matrices to the video latent in a pixel-wise manner.
@@ 4.1 Experiment Settings
Vision-Language Models. We pair our pipeline with four representative VLMs: the closed-source GPT-4o and GPT-4.1 (strong general-purpose multimodal baselines), InternVL3-14B (one of the most capable open-source VLMs to date), and o1, a reinforcement-learning–fine-tuned model that performs test-time chain-of-thought scaling. This mix spans both proprietary and fully open ecosystems as well as models with and without explicit reasoning scaffolds.
World Models. Our experiments use two distinct video world models. (i) Search World Model (SWM), the world model we trained, introduced in Sec. 3.4; and (ii) Stable-Virtual-Camera (SVC), a recently released, publicly available generator that produces geometrically stable novel views.
@@ 4.2 Experiments Results
The SAT-Real split comprises 150 real-image queries spanning indoor and outdoor scenes. Table 1 shows that augmenting each VLM with MindJourney results in uniform and sizeable gains. Average top-1 accuracy rises by 7.7 percentage points, and GPT-4o achieves the largest single boost at over 10%. Remarkably, GPT-4.1 with our method already surpasses vanilla o1; when o1 itself is paired with MindJourney it sets a new state of the art on SAT-Real. These findings confirm that world-model-driven test-time scaling complements RL-based scaling and our method generalizes to real, outdoor scenarios.
@@
Because the synthetic split contains 4 000 questions, we evaluate on a random 500-question subset to keep the o1 runs tractable. Results in Table 2 reproduce the pattern seen on real images: mean accuracy improves by 8.0% in average. Again, the performance of GPT-4.1 with our method surpasses vanilla o1, and o1 with our method improves its performance. Moreover, across all five SAT question types the highest score is always achieved by a MindJourney-augmented model, underscoring that the proposed framework consistently improves reasoning across all question types.
@@ 4.3 Ablation Study
In our ablation study, we ablate the hyperparameters of our search method. Specifically, for gpt-4o augmented with SWM, we ablate with search step $n\in\{1,2,3\}$ and VLM pruning threshold $\gamma\in\{4,6,8\}$. We experimented on these values on both SAT-Real and SAT-Synthesized. Accuracy is reported on both SAT-Real and the 500-question SAT-Synthesized subset, allowing us to evaluate how each setting affects performance in real-image and synthetic scenarios.
Search Depth. Panel (a) of Fig. 4 shows that accuracy for threshold 4 and 6 on SAT-Real peaks at search steps 2 and then drops slightly at step 3. The decline stems from the limits of our world model, SWM: trained predominantly on indoor synthetic data, it can struggle to faithfully simulate outdoor views once the imagined trajectory strays too far from the reference frame. On SAT-Synthesized—whose scenes are closer to the SWM’s training distribution—performance continues to rise through step 3 for both threshold 8 and 6, confirming that deeper exploration remains beneficial when the world model can faithfully render long roll-outs. More visual results can be found in the Appendix.
Pruning Threshold. Panels (a) and (b) also show the importance of the VLM score threshold. A lenient threshold lets many low-quality views into the evidence buffer, diluting the signal and lowering accuracy. The effect is stronger on the Real split, where generation quality is lower, so the threshold has a larger impact there.
@@
Figure 4: Inference Steps vs. Accuracy. Accuracy on SAT-Real and SAT-Synthesized with different VLM Thresholds and inference steps.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
We are grateful to Anushka Agarwal for assistance with the baseline code, and to Jiaben Chen, Zeyuan Yang, Lixing Fang, Haoyu Zhen, and many other friends for their helpful feedback and insightful discussions.
@@ A.1.1 Search Configurations
We use the same search configuration for every experiment: a search depth of $n=3$ steps, a beam size of 2, exploration and helpfulness thresholds $\gamma_{\text{exp}}=8$ and $\gamma_{\text{help}}=8$, and a maximum trajectory length of 8 starting from the given reference image. During each expansion, we allow up to $k=3$ consecutive repetitions per primitive action; each forward step moves the agent 0.250.25m and each rotation step turns it by $9^{\circ}$. We prepared more visual examples about the Trajectory Expansion process under our experiment settings at Fig. 5 and Fig. 6
@@ A.1.2 Computational Resources
All inference experiments were run on high‑performance NVIDIA GPUs: when using Search World Model as the world model, we employed A40 GPUs with 40GB of VRAM; when using Stable‑Virtual‑Camera as the world model, we ran on H100 GPUs with 80GB of VRAM; and for all experiments combining the InternVL3‑14B VLM with the Search World Model, we also used H100 GPUs to accommodate the larger memory footprint of the vision–language model.
@@ A.2.1 Dataset
The training set for our Search World Model (SWM) comprises three components: HM3D, DL3DV‑10K, and RealEstate10K (Szot et al., 2022; Ling et al., 2024; Zhou et al., 2018).
HM3D. We generate 50K simulated navigation clips using Habitat on HM3D scenes. For each episode, we sample a random start and goal position and follow the shortest path for up to 500 steps. To improve robustness to camera tilt, we uniformly draw an initial pitch from $30^{\circ},20^{\circ}]$, and in 10% of episodes we hold the agent at a fixed location and vary only its pitch by rotating up and down.
@@
DL3DV‑10K. Similarly, we incorporate 10K real outdoor videos from DL3DV‑10K to capture natural scenes and broaden our model’s generalization to exterior settings.
@@ A.2.2 Implementation Details
We adopt Wan2.2-TI2V‑5B (Wan et al., 2025) as our backbone. We followed the implementation of ReCamMaster, which conditions video generation on the target camera poses by encoding each frame’s extrinsic matrix (3×4 rotation–translation) with a small learnable camera encoder and adding this signal to the visual features inside every diffusion-transformer block, right after spatial attention and before 3D (spatio-temporal) attention. During the training, we only finetuned the camera encoder and a newly introduced projector. For more details, please refer to ReCamMaster’s open-sourced codebase.
@@ A.2.3 Training Details
We subsample training clips with a frame‐skip stride uniformly drawn between 1 and 3 to expose the model to varied camera motions. Optimization is performed with Adam and a linear warmup schedule to a peak learning rate of 3e53e-5, using bfloat16 precision for efficiency and clipping gradients to a maximum norm of 1.0 for stability. All video‐diffusion models are trained on eight NVIDIA H100 GPUs over approximately three days.
@@ B.1.1 Case 1: Inaccurate Forward Movement
In group (a), the imagined trajectories systematically under‑ or over‑estimate forward translations: the actual step lengths no longer match the intended distances, and the displacement between successive frames varies unpredictably. As a result, the agent repeatedly overshoots its targets and exhibits jittery, erratic motion within the simulated environment.
@@ C.2 Ablation on Trajectoy Description
C.3 Ablation on Time Consumption
To assess the time efficiency of our method, we evaluated it with GPT-4o as the VLM while varying the number of inference steps. We also reduced the SWM’s video-generation iterations from 50 to 20 to examine the trade-off between runtime and accuracy. The results show that fewer SWM iterations markedly speed up inference but degrade the quality of the generated video, which in turn lowers accuracy. Conversely, increasing the number of inference steps forces the SWM to predict more frames and provides more images to the VLM, thereby increasing both the SWM time and the VLM search/QA times.
However, we also observe that our method does not degrade significantly with 2 search steps and 20 iterations. Therefore, while our current setting optimizes for performance, step 2 and iter 20 is a much more cost-effective and recommended setting practically.
Table 4: SWM/VLM Time Consumption and Accuracy
Setting SWM time(s) Search time(s) Q&A time(s) Total(s) ACC(%)
step:1 iter:20 11.49 3.04 1.74 16.27 64.7
step:2 iter:20 19.68 5.53 1.71 26.93 68.6
step:3 iter:20 29.37 8.00 3.24 40.60 69.3
step:1 iter:50 30.57 2.94 1.84 35.36 65.3
step:2 iter:50 63.62 5.52 1.79 70.93 69.3
step:3 iter:50 149.75 8.30 3.43 161.48 70.6